# PLD: A Choice-Theoretic List-Wise Knowledge Distillation

### v2

@@ Abstract
Knowledge distillation is a model compression technique in which a compact "student" network is trained to replicate the predictive behavior of a larger "teacher" network. In logit-based knowledge distillation, it has become the de facto approach to augment cross-entropy with a distillation term. Typically, this term is either a KL divergence-matching marginal probabilities or a correlation-based loss-capturing intra and inter-class relationships-but in every case it sits as an "add-on" to cross-entropy, with its own weight that must be carefully tuned. In this paper, we adopt a choice-theoretic perspective and recast knowledge distillation under the Plackett-Luce model by interpreting teacher logits as "worth" scores. We introduce Plackett-Luce Distillation (PLD), a weighted list-wise ranking loss in which the teacher model transfers knowledge of its full ranking of classes, weighting each ranked choice by its own confidence. PLD directly optimizes a single "teacher-optimal" ranking-true label first, followed by the remaining classes in descending teacher confidence-yielding a convex, and translation-invariant surrogate that subsumes weighted cross-entropy. Empirically, on standard image-classification benchmarks, PLD improves Top-1 accuracy by an average of +0.42% over (DIST)huang2022knowledge and +1.04% over KDhinton2015distilling in homogeneous settings, and by +0.48% and +1.09% over (DIST) and KD, respectively, in heterogeneous settings.
@@ 1 Introduction
Deep neural networks (DNNs) have achieved remarkable success across a wide array of tasks-from image classification and object detection to semantic segmentation and beyond yang2022cloud ; wang2024towards . Accuracy, generalization, and robustness typically improve with increasing depth and parameter count wang2024end , but these gains come at the cost of significantly higher computational and memory demands, limiting deployment in resource-constrained environments such as mobile and embedded devices.
Knowledge distillation (KD) hinton2015distilling addresses this by training a compact “student” network to mimic a large “teacher” network. Classically, KD minimizes a weighted sum of cross-entropy on hard labels and a temperature-scaled Kullback-Leibler (KL) divergence between teacher and student logits. This simple framework enriches the student’s learning signal and has seen widespread adoption for model compression and robustness enhancement gou2021knowledge ; song2022spot .
Despite its popularity, recent studies reveal a paradox: distilling from ever-larger or more accurate teachers can sometimes degrade student performance, particularly under capacity mismatch mirzadeh2020improved ; son2021densely ; huang2022knowledge . A variety of techniques have been proposed to address this issue-teacher-assistant frameworks mirzadeh2020improved ; son2021densely , selective distillation cho2019efficacy ; zhu2022teach , and auxiliary classifiers liang2024neighbor -but these approaches introduce additional architectural components or training stages. An alternative line of work, DIST huang2022knowledge , remains within the classical KD pipeline yet replaces the KL-based distillation term. DIST argues that matching only marginal probabilities via KL divergence fails to preserve the relational structure encoded in the teacher’s outputs. Instead, DIST employs a Pearson correlation-based loss to preserve both inter-class correlations within each prediction and intra-class correlations across examples. Nevertheless, like other distillation methods, DIST still requires a separate cross-entropy term. As Figure 1(a) shows, reducing the weight on this cross-entropy term can improve accuracy, but removing it entirely leads to performance drop. Consequently, a unified objective that subsumes both cross-entropy and distillation-deriving its weights from the teacher’s confidences rather than manual tuning-would be highly desirable.
@@
To overcome these limitations, we propose a list-wise, choice-theoretic perspective on distillation. We interpret logits as “worth” scores in the classical Plackett-Luce permutation model luce1959individual ; plackett1975analysis , derive a single teacher-optimal ranking-true label first, followed by the teacher’s remaining classes in descending confidence-and impose it on the student via the unsoftened PL likelihood. The PL model is grounded in Luce’s Choice Axiom, which ensures that choice-probability ratios are invariant to irrelevant alternatives, and in Plackett’s extension to full rankings, which yields a coherent distribution over permutations by sequential removal. Because the PL model’s first selection term coincides with cross-entropy on the true label, it inherently subsumes the standard KD cross-entropy loss. However, similar to reddi2021rankdistil ; lan2014position , we observe that proper weighting of each subsequent selection is crucial (see Sec. 5). By weighting step $k$ by the teacher’s softmax mass $\alpha_{k}$, we introduce the Plackett-Luce Distillation (PLD) Loss, a convex, translation-invariant surrogate that unifies cross-entropy, ListMLE xia2008listwise , and P-ListMLE lan2014position as special cases. This choice-theoretic foundation captures the full ranking structure provided by the teacher and enables efficient, gradient-based optimization, eliminating the need for separate cross-entropy terms or manual weight adjustments.
@@
  • We empirically demonstrate on the image-classification benchmark that PLD consistently outperforms both traditional KD hinton2015distilling and DIST huang2022knowledge : in homogeneous settings, PLD improves Top-1 accuracy by an average of +0.42% over DIST and +1.04% over KD; with a fixed ResNet-50 student distilled from diverse teachers, it achieves gains of +0.48% versus DIST and +1.09% versus KD.
@@ 2.1 Knowledge Distillation and Capacity-Mismatch Remedies
Knowledge distillation hinton2015distilling , inspired by model-bootstrapping techniques zhou2016learnware , trains a compact student to mimic a larger teacher by minimizing a temperature-scaled KL divergence between their outputs logits (softened probabilities) to transfer accuracy and robustness gou2021knowledge while achieving model compression song2022spot . Over time, this simple yet effective framework has become a cornerstone for compressing and enhancing deep neural networks. However, a range of studies have shown that students distilled from very large or highly accurate teachers can underperform, even under adversarially robust settings cho2019efficacy ; park2019relational ; mirzadeh2020improved ; son2021densely ; zhu2021student ; wang2022efficient ; zhu2022teach ; huang2022knowledge ; rao2023parameter ; liang2024neighbor ; yuan2024student ; yin2024adversarial . This counterintuitive degradation is mainly ascribed to a capacity mismatch between teacher and student. To address this, architecture-level innovations introduce intermediate or auxiliary models-Teacher Assistant KD (TAKD) employs a mid-sized assistant to smooth the gap mirzadeh2020improved , Densely Guided KD (DGKD) aggregates multiple assistants for richer supervision son2021densely , Neighbor Self-KD (NSKD) adds auxiliary classifiers within student layers liang2024neighbor , and Student Customized KD (SCKD) gates the distillation loss based on gradient alignment with the student’s own objective zhu2021student . Complementarily, teacher-knowledge regularization techniques refine the source signals: early-stopped teachers often outperform fully converged ones cho2019efficacy , CheckpointKD selects intermediate checkpoints to avoid over-specialization wang2022efficient , undistillable classes are excluded to focus supervision on reliable predictions zhu2022teach , adapter modules control label smoothness rao2023parameter , and Student-friendly KD (SKD) simplifies teacher outputs through a softening-simplifier pipeline yuan2024student .
@@ 2.2 Logit-Based Distillation Beyond KL Divergence
This trend is reflected in recent logit-based distillation methods that seek alternatives to KL divergence. Relational KD (RKD) park2019relational transfers mutual relations among data examples via distance and angle-wise losses that penalize structural discrepancies. DIST huang2022knowledge introduced a Pearson correlation coefficient-based loss to explicitly capture the teacher’s intrinsic inter and intra-class relations. Correlation Matching KD (CMKD) niu2024efficient reinterprets inter-class rankings in terms of the decision boundary and employs both Pearson and Spearman correlation losses, dynamically weighted by sample difficulty using a differentiable sorting operation. Additionally, it was observed fan2024revisit that better teacher calibration correlates with improved KD performance under standard methods and therefore was advocated for calibration-insensitive metrics such as ranking-based losses.These methods are computationally efficient-requiring no complex training pipelines or architectural overhead-yet achieve competitive performance. This observation motivates our search for a similarly efficient, calibration-insensitive objective grounded in list-wise ranking theory.
@@ 2.3 Ranking-Based and List-Wise Losses for Distillation
Early ranking methods focused on pairwise losses, where each pair of items is classified as correctly or incorrectly ordered-e.g., the ranking SVM (herbrich1999support, ; joachims2002optimizing, ). Listwise approaches, which operate over entire item lists, later gained traction in information retrieval and machine learning (cao2007learning, ; xia2008listwise, ). Many draw on classical statistical ranking models-most notably Luce’s choice model (luce1959individual, ) and the Plackett-Luce permutation distribution (plackett1975analysis, ). However, as the number of items grows, even stochastic top-$k$ extensions like ListNet (luo2015stochastic, ) become impractical, and these losses-designed for direct supervision-do not naturally extend to knowledge distillation.
More recent work applies ranking losses to classification and distillation. For multiclass tasks, listwise losses have been proposed wang2021rank4class ; frydenlund2022language , while distillation under a position-aware binary ranking loss has been studied tang2018ranking ; gao2020understanding . The method in tang2018ranking discards items that the teacher ranks low, whereas lee2019collaborative adapts ranking-based distillation to collaborative filtering. Similarly, RankDistil reddi2021rankdistil preserves the teacher’s top-$k$ ordering by matching the student’s item order to the teacher’s and penalizing lower-ranked items.
Building on ListMLE xia2008listwise , we directly optimize the likelihood of a single teacher-optimal permutation under the Plackett-Luce model. Inspired by RankDistil reddi2021rankdistil and position-aware ListMLE (P-ListMLE) lan2014position , we assign greater weight to top-ranked classes than to those ranked lower. Unlike RankDistil, which is primarily designed for general ranking scenarios and optimizes top-$k$ metrics, our focus is on classification tasks where top-1 accuracy is of primary importance.
@@ 3.2 Plackett-Luce (PL) Ranking Model
Let $s\in\mathbb{R}^{C}$ be the logit vector produced by the network. We write $s_{i}$ for its $i$th component and interpret $i\succ j\quad\Longleftrightarrow\quad s_{i}>s_{j}$ as “class $i$ preferred to class $j$.” A full ranking is a permutation $\pi=(\pi_{1},\dots,\pi_{C})\in S_{C}$. The Plackett-Luce model assigns each class $i$ a strictly positive worth $w_{i}=\phi(s_{i})>0,$ where $\phi:\mathbb{R}\to\mathbb{R}_{>0}$ is strictly increasing (we take $\phi(s)=e^{s}$). It defines a probability over permutations by
@@
Since PL factorizes the full $C!$-way ranking distribution into chained softmax terms, it provides an efficient likelihood for any fixed permutation (xia2008listwise, ). In Section 4 we fix a “teacher-optimal” permutation $\pi^{*}$ and minimize its negative log-likelihood under the student, thereby enforcing both top-1 correctness and richer inter-class ordering information.
@@ 4 Method
We derive our PLD objective in two stages: First, we extract a teacher-optimal permutation from the teacher’s logits via the Plackett-Luce model (Sec. 4.1), thereby identifying the core ranking target for distillation. Second, we introduce a confidence-weighted likelihood to obtain the final Plackett-Luce Distillation (PLD) loss (Sec. 4.2).
@@ 4.1 Teacher-Optimal Permutation under Plackett-Luce
From the Plackett-Luce ranking viewpoint, a classifier’s logits define a sequential choice process over the $C$ classes. In the first selection step, the model chooses one class out of all $C$ in proportion to $\exp(s_{i})$, exactly matching the softmax probability used in standard cross-entropy. Consequently, minimizing cross-entropy enforces only this first-choice probability-i.e., the probability of picking the correct class first-while remaining agnostic to the ordering of the other $C-1$ classes. In contrast, the full Plackett-Luce model assigns distinct probabilities to all $C!$ permutations by chaining these softmax-style choices at each subsequent step. Thus, cross-entropy can be seen as a relaxed Plackett-Luce loss that cares only about the top-1 selection.
If we augment cross-entropy with a knowledge-distillation term-whether via KL matching of softened distributions or via correlation-based losses-the resulting objective still places its primary emphasis on the true class (via the cross-entropy term) while the distillation term modulates the student’s remaining logits. Viewed through the Plackett-Luce lens, this corresponds to enforcing the first selection to be the ground-truth label, exactly as in cross-entropy, and then shaping the subsequent choices according to the teacher’s preferences rather than forcing an exact match of the full distribution. Concretely, since Plackett-Luce makes each selection independently-choosing $\pi_{k}$ among the remaining classes in proportion to their worth-we fix the first pick to the true label and let the teacher’s descending-logit ordering guide all later picks. Denoting the teacher’s logit vector by $t=(t_{1},\dots,t_{C})\in\mathbb{R}^{C}$, we refer to the resulting fixed ordering
@@
It was shown in xia2008listwise that one can construct surrogate losses on a single Plackett-Luce permutation-using, for example, likelihood (ListMLE), cosine similarity, or cross-entropy-that are all amenable to gradient-based optimization. Among these, the maximum-likelihood (ListMLE) surrogate stands out for its simplicity and for satisfying desirable properties such as consistency, soundness, continuity, differentiability, and convexity, while also exhibiting strong empirical performance. Accordingly, we adopt the ListMLE approach: we place an empirical one-hot target on the teacher-optimal permutation $\pi^{*}$ and define our unweighted loss as its negative log-likelihood under the student’s logits,

### v3

@@ Abstract
Knowledge distillation is a model compression technique in which a compact "student" network is trained to replicate the predictive behavior of a larger "teacher" network. In logit-based knowledge distillation, it has become the de facto approach to augment cross-entropy with a distillation term. Typically, this term is either a KL divergence that matches marginal probabilities or a correlation-based loss that captures intra- and inter-class relationships. In every case, it acts as an additional term to cross-entropy. This term has its own weight, which must be carefully tuned. In this paper, we adopt a choice-theoretic perspective and recast knowledge distillation under the Plackett–Luce model by interpreting teacher logits as "worth" scores. We introduce Plackett-Luce Distillation (PLD), a weighted list-wise ranking loss. In PLD, the teacher model transfers knowledge of its full ranking of classes, weighting each ranked choice by its own confidence. PLD directly optimizes a single "teacher-optimal" ranking. The true label is placed first, followed by the remaining classes in descending teacher confidence. This process yields a convex and translation-invariant surrogate that subsumes weighted cross-entropy. Empirically, across CIFAR-100, ImageNet-1K, and MS-COCO, PLD achieves consistent gains across diverse architectures and distillation objectives, including divergence-based, correlation-based, and feature-based methods, in both homogeneous and heterogeneous teacher–student pairs.
@@ 1 Introduction
Deep neural networks (DNNs) have achieved remarkable success across a wide array of tasks-from image classification and object detection to semantic segmentation and beyond yang2022cloud ; wang2024towards . Accuracy, generalization, and robustness usually improve as models become deeper and larger wang2024end . However, these gains require more computation and memory, which limits deployment on mobile or embedded devices.
Knowledge distillation (KD) hinton2015distilling addresses this by training a compact "student" network to mimic a large "teacher" network. Classically, KD minimizes a weighted sum of cross-entropy on hard labels and a temperature-scaled Kullback-Leibler (KL) divergence between teacher and student logits. This simple framework enriches the student’s learning signal and has seen widespread adoption for model compression and robustness enhancement gou2021knowledge ; song2022spot .
Despite its popularity, recent studies reveal a paradox. Distilling from larger or more accurate teachers can sometimes degrade the student’s performance, especially when there is a capacity mismatch mirzadeh2020improved ; son2021densely ; huang2022knowledge . Several methods have been proposed to address this issue. These include teacher-assistant frameworks mirzadeh2020improved ; son2021densely , selective distillation cho2019efficacy ; zhu2022teach , and auxiliary classifiers liang2024neighbor . However, these methods often add extra architectural components or training stages. An alternative line of work, DIST huang2022knowledge , remains within the classical KD pipeline yet replaces the KL-based distillation term. DIST argues that matching only marginal probabilities via KL divergence fails to preserve the relational structure encoded in the teacher’s outputs.
Instead, DIST employs a Pearson correlation-based loss to preserve both inter-class correlations within each prediction and intra-class correlations across examples. Nevertheless, like other distillation methods, DIST still requires a separate cross-entropy term. As Figure 1(a) shows, reducing the weight on this cross-entropy term can improve accuracy, but removing it entirely leads to performance drop. Therefore, a unified objective that combines cross-entropy and distillation is desirable. Such an objective should derive its weights from the teacher’s confidence rather than manual tuning.
@@
To overcome these limitations, we propose a list-wise, choice-theoretic perspective on distillation. We interpret logits as "worth" scores under the classical Plackett–Luce model luce1959individual ; plackett1975analysis . We then derive a single teacher-optimal ranking, where the true label comes first and the remaining classes follow in descending teacher confidence. This ranking is imposed on the student using the unsoftened Plackett–Luce likelihood. The PL model is based on Luce’s Choice Axiom, which ensures that choice probabilities are invariant to irrelevant alternatives. Plackett’s extension generalizes this to full rankings, forming a coherent distribution over permutations through sequential removal. The first selection term of the PL model matches the cross-entropy on the true label. Thus, it naturally includes the standard KD cross-entropy loss. However, similar to reddi2021rankdistil ; lan2014position , we observe that proper weighting of each subsequent selection is crucial (see Sec. 5). We weight each step $k$ by the teacher’s softmax mass $\alpha_{k}$. This leads to the Plackett–Luce Distillation (PLD) Loss, a convex and translation-invariant surrogate. PLD unifies cross-entropy, ListMLE xia2008listwise , and P-ListMLE lan2014position as special cases. This choice-theoretic foundation captures the full ranking structure provided by the teacher and enables efficient, gradient-based optimization, eliminating the need for separate cross-entropy terms or manual weight adjustments.
@@
  • We empirically demonstrate on CIFAR-100, ImageNet-1K, and MS-COCO that PLD achieves consistent gains across diverse architectures and distillation objectives, including divergence-based, correlation-based, and feature-based methods, in both homogeneous and heterogeneous teacher–student pairs.
@@ 2.1 Knowledge Distillation and Capacity-Mismatch Remedies
Knowledge distillation (KD) hinton2015distilling , inspired by model-bootstrapping techniques zhou2016learnware , trains a compact student to mimic a larger teacher. It minimizes a temperature-scaled KL divergence between their output logits (softened probabilities). This process transfers both accuracy and robustness gou2021knowledge while enabling model compression song2022spot . Over time, this simple yet effective framework has become a cornerstone for compressing and enhancing deep neural networks. However, a range of studies have shown that students distilled from very large or highly accurate teachers can underperform, even under adversarially robust settings cho2019efficacy ; park2019relational ; mirzadeh2020improved ; son2021densely ; zhu2021student ; wang2022efficient ; zhu2022teach ; huang2022knowledge ; rao2023parameter ; liang2024neighbor ; yuan2024student ; yin2024adversarial . This counterintuitive degradation is mainly ascribed to a capacity mismatch between teacher and student. To address this issue, several architecture-level methods introduce intermediate or auxiliary models. Teacher Assistant KD (TAKD) employs a mid-sized assistant to bridge the gap mirzadeh2020improved . Densely Guided KD (DGKD) aggregates multiple assistants for richer supervision son2021densely . Neighbor Self-KD (NSKD) adds auxiliary classifiers within student layers liang2024neighbor . Student Customized KD (SCKD) adjusts the distillation loss based on gradient alignment with the student’s objective zhu2021student . Teacher-knowledge regularization techniques also refine the teacher’s signal. Early-stopped teachers often outperform fully converged ones cho2019efficacy . CheckpointKD selects intermediate checkpoints to prevent over-specialization wang2022efficient . Some methods exclude undistillable classes to focus supervision on reliable predictions zhu2022teach . Adapter modules control label smoothness rao2023parameter , and Student-friendly KD (SKD) simplifies teacher outputs through a softening-simplifier pipeline yuan2024student .
@@ 2.2 Logit-Based Distillation Beyond KL Divergence
This trend is reflected in recent logit-based distillation methods that seek alternatives to KL divergence. Relational KD (RKD) park2019relational transfers mutual relations among data examples via distance and angle-wise losses that penalize structural discrepancies. DIST huang2022knowledge introduced a Pearson correlation coefficient-based loss to explicitly capture the teacher’s intrinsic inter and intra-class relations. Correlation Matching KD (CMKD) niu2024efficient reinterprets inter-class rankings in terms of the decision boundary and employs both Pearson and Spearman correlation losses, dynamically weighted by sample difficulty using a differentiable sorting operation. It was also observed fan2024revisit that better teacher calibration correlates with improved KD performance under standard methods. This finding motivated calibration-insensitive metrics such as ranking-based losses. These methods are computationally efficient, require no complex training pipelines or architectural overhead, yet achieve competitive performance. This observation motivates our search for a similarly efficient, calibration-insensitive objective grounded in list-wise ranking theory.
@@ 2.3 Ranking-Based and List-Wise Losses for Distillation
Early ranking methods focused on pairwise losses. In these methods, each pair of items is classified as correctly or incorrectly ordered, as in the ranking SVM (herbrich1999support, ; joachims2002optimizing, ). Listwise approaches, which operate over entire item lists, later gained traction in information retrieval and machine learning (cao2007learning, ; xia2008listwise, ). Many draw on classical statistical ranking models-most notably Luce’s choice model (luce1959individual, ) and the Plackett-Luce permutation distribution (plackett1975analysis, ). As the number of items increases, even stochastic top-$k$ extensions such as ListNet (luo2015stochastic, ) become impractical. Moreover, these losses are designed for direct supervision and do not naturally extend to knowledge distillation.
More recent work applies ranking losses to classification and distillation. For multiclass tasks, listwise losses have been proposed wang2021rank4class ; frydenlund2022language , while distillation under a position-aware binary ranking loss has been studied tang2018ranking ; gao2020understanding . The method in tang2018ranking removes items that the teacher ranks low. In contrast, lee2019collaborative adapts ranking-based distillation to collaborative filtering. Similarly, RankDistil reddi2021rankdistil preserves the teacher’s top-$k$ ordering by matching the student’s item order to the teacher’s and penalizing lower-ranked items.
Building on ListMLE xia2008listwise , we directly optimize the likelihood of a single teacher-optimal permutation under the Plackett-Luce model. Inspired by RankDistil reddi2021rankdistil and position-aware ListMLE (P-ListMLE) lan2014position , we assign greater weight to top-ranked classes than to those ranked lower. Unlike RankDistil, which is designed for general ranking tasks and optimizes top-$k$ metrics, our method focuses on classification tasks. In this setting, top-1 accuracy is most important.
@@ 3.2 Plackett-Luce (PL) Ranking Model
Let $s\in\mathbb{R}^{C}$ be the logit vector produced by the network. We write $s_{i}$ for its $i$th component and interpret $i\succ j\quad\Longleftrightarrow\quad s_{i}>s_{j}$ as "class $i$ preferred to class jj." A full ranking is a permutation $\pi=(\pi_{1},\dots,\pi_{C})\in S_{C}$. The Plackett-Luce model assigns each class $i$ a strictly positive worth $w_{i}=\phi(s_{i})>0,$ where $\phi:\mathbb{R}\to\mathbb{R}_{>0}$ is strictly increasing (we take $\phi(s)=e^{s}$). It defines a probability over permutations by
@@
The PL model factorizes the full $C!$-way ranking distribution into a chain of softmax terms. This factorization provides an efficient likelihood for any fixed permutation (xia2008listwise, ). In Section 4, we fix a "teacher-optimal" permutation $\pi^{*}$ and minimize its negative log-likelihood under the student. This approach enforces both top-1 correctness and richer inter-class ordering information.
@@ 4 Method
We derive the PLD objective in two stages. First, we extract a teacher-optimal permutation from the teacher’s logits using the Plackett–Luce model (Sec. 4.1). This step identifies the ranking target for distillation. Second, we introduce a confidence-weighted likelihood to form the final Plackett–Luce Distillation (PLD) loss (Sec. 4.2).
@@ 4.1 Teacher-Optimal Permutation under Plackett-Luce
From the Plackett-Luce ranking viewpoint, a classifier’s logits define a sequential choice process among the $C$ classes. In the first selection step, the model chooses one class out of all $C$ in proportion to $\exp(s_{i})$, exactly matching the softmax probability used in standard cross-entropy. Consequently, minimizing cross-entropy enforces only the first-choice probability, i.e., the probability of selecting the correct class first. It remains agnostic to the ordering of the remaining C1C-1 classes. In contrast, the full Plackett–Luce model assigns probabilities to all $C!$ possible permutations. It does this by chaining softmax-style selections at each step. Thus, cross-entropy can be seen as a relaxed Plackett-Luce loss that cares only about the top-1 selection.
If we augment cross-entropy with a knowledge-distillation term, such as KL matching or correlation-based loss, the objective still focuses mainly on the true class through the cross-entropy component. The distillation term only adjusts the remaining logits. Viewed through the Plackett–Luce perspective, this process enforces the first selection to be the ground-truth label, as in cross-entropy. It then shapes later choices according to the teacher’s preferences instead of matching the entire distribution. Concretely, since Plackett-Luce makes each selection independently-choosing $\pi_{k}$ among the remaining classes in proportion to their worth-we fix the first pick to the true label and let the teacher’s descending-logit ordering guide all later picks. Denoting the teacher’s logit vector by $t=(t_{1},\dots,t_{C})\in\mathbb{R}^{C}$, we refer to the resulting fixed ordering
@@
As shown in xia2008listwise , one can construct surrogate losses on a single Plackett–Luce permutation. Examples include likelihood-based (ListMLE), cosine-similarity, and cross-entropy variants, all suitable for gradient-based optimization. Among these, the likelihood (ListMLE) surrogate stands out for its simplicity and for satisfying desirable properties such as consistency, soundness, continuity, differentiability, and convexity, while also exhibiting strong empirical performance. We adopt the ListMLE approach. We place an empirical one-hot target on the teacher-optimal permutation $\pi^{*}$ and define its negative log-likelihood under the student’s logits as the unweighted loss.